A rectangle has a length of 3x +9 and a width of 5x -4. The perimeter of the rectangle is 106
units. Find the width of the rectangle.

Found 2 solutions by stanbon, macston:
Answer by stanbon(75887)

A rectangle has a length of 3x +9 and a width of 5x -4. The perimeter of the rectangle is 106
units. Find the width of the rectangle.
------
Length + width = 53
3x+9 + 5x-4 = 8x+5 = 53
8x = 48
x = 6
-----
Width = 5x-4 = 5*6-4 = 26
-----
Cheers,
Stan H.
----------





Answer by macston(5194)

L=length=3x+9; W=width=5x-4; P=perimeter=106
P=2(L+W)
106=2(3x+9+5x-4) Divide each side by 2
53=8x+5 Subtract 5 from each side.
48=8x Divide each side by 8.
6=x
W=5x-4=5(6)-4=30-4=26 ANSWER: The width is 26 units.
CHECK:
L=3x+9=3(6)+9=18+9=27 The length is 27 units
P=2(L+W)
106=2(27+26)
106=2(53)
106=106